james and thomas are in two cities which are 186 miles apart and travel toward each other, james, average rate was 32 mph and thomas average rate was 36mph if james started at 9:00 A.M and thomas started at 9:30 A.M, at what time did they meet?
Answer by stanbon(75887)

james and thomas are in two cities which are 186 miles apart and travel toward each other, james, average rate was 32 mph and thomas average rate was 36mph if james started at 9:00 A.M and thomas started at 9:30 A.M, at what time did they meet?
------------------
James DATA:
Rate = 32 mph ; time = x hrs ; distance = 12x miles
---------------------
Thomas DATA:
Rate = 36 mph ; tune = (x-(1/2)) hrs; distance = 36(x-1/2)= 36x-18
-----------------------
EQUATION:
distance + distance = 186
12x + 36x-18 = 186
48x = 204
x = 4.25 hours.
--------------------
Meeting time = 9:00 AM + 4 1/4 hrs = 1:15 PM
=====================
Cheers
Stan H.